OpenStudy (anonymous):

The perimeter of a square is directly proportional to the length of one of its sides. The perimeter is 28 when the length of a side is 7. What is the constant of proportionality? 4 14 196

jimthompson5910 (jim_thompson5910):

let k = constant of proportionality P = k*s 28 = k*7 k = 28/7 k = ???

OpenStudy (anonymous):

4 haha I thought so (: thanks

jimthompson5910 (jim_thompson5910):

yw